Operation Summary for June 21, 2010 - June 23, 2010
Monday June 21
The day shift began with Operations monitoring store 7908 (90.28E30) and the Recycler stash (354.4E10), with the Antiproton Source (Pbar) stacking (18.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson FTTB T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, Operations terminated store 7908 at 10:05 AM, put the TeV into HEP shot setup, and established store 7909 at 11:30 AM with an average initial luminosity (AIL) of 364E30. TeV personnel conducted a short round of beam studies during shot setup.
During the evening shift, Controls personnel swapped out a microprocessor card for the AP60 BPM VME crate. The H- Source cesium tube temperature increased rapidly causing the cathode and source temperatures to rise. Operators accessed the dome to decrease the temperature and stabilize the source. The temperature had begun decreasing by the end of shift.
Tuesday June 22
The midnight shift began with Operations monitoring store 7909 (83E30) and the stash (315E10), with Pbar stacking (18E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 7909 at 3 AM, put the TeV into shot setup, and established store 7910 with an AIL of 302E30; the AIL was lower than expected due to an antiproton emittance blow up. Operators monitored the H- Source temperature, which was decreasing very slowly. Linac personnel arrived near the end of the shift to investigate further.
During the day shift, Operations reported 30 minutes of downtime due to Preaccelerator personnel accessing the H- Source dome and to Linac personnel repairing a RF station (LRF3). Operations established beam to T-977.
During the evening shift, Operations disabled the T-977 beam. Operations terminated store 7910 at 8:04 PM, put the TeV into shot setup at 8:15 PM, and established store 7911 at 9:30 PM with an AIL of 318E30. Antiproton emittances were higher than normal; operators adjusted the tune.
Wednesday June 23
The midnight shift began with Operations monitoring store 7911 (182.38E30) and the stash (112.4E10), with Pbar stacking (10.3E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to T-977. Operations disabled the NuMI beam at 7:15 AM to allow for horn power supply repairs.
The Plans for Wednesday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 06/23/2010 email Fermilab|
|Security, Privacy, Legal|